OpenStudy (anonymous):

What is the y-value of cos(x) when x = -360°? A. -1 B. 0 C. 1/2 D. 1

OpenStudy (lgbasallote):

\[y = \cos x\] \[y = \cos (-360)\] use your calculator to solve...

OpenStudy (anonymous):

is it 1

hartnn (hartnn):

yup, its 1 :)

OpenStudy (anonymous):

thanx

OpenStudy (raden):

don't forget the property of trygonometri one of that : cos(-x) = cos(x)